OpenStudy (ineedhelpnow):

solve using substitution: 2x+3y=11 x-4y=0

OpenStudy (anonymous):

x-4y=0 add 4y to both side

OpenStudy (anonymous):

x=4y

OpenStudy (anonymous):

2(4y)+3y=11

OpenStudy (anonymous):

8y+3y=11

OpenStudy (anonymous):

11y=11

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

x=4y substitute x in 1st equation to get y value then substitute y value in one of the equation to get x value

OpenStudy (anonymous):

2x + 3y = 11 transpose 3y to right hand side 2x = 11-3y divide both sides by 2 2x/2 = ( 11-3y ) /2 x = (11-3y)/2 put the value of x in the second equation (11-3y)/2 - 4y = 0 (11-3y-8y)/2 = 0 (11-11y) = 0*2 11-11y = 0 transpose 11 to right hand side -11y =0-11 -11y = -11 divide both sides by -11 y = 1